If a+b+c=0=p+q+r then the value of determinant              pa    qb    rc                                                                                                                                                        qc   ra     pb                                                                                                                                                                    rb    pc     qa          is? how

Grade:12

1 Answers

APURV GOEL
12 years ago

When we open the matrice we get

              prq(a3 + b3 + c3) - abc(p3 + q3 +r3)

    we have a+b+c = p+q+r =0

     so a3+b3+c3 = 3abc    and     p3+r3+q3 = 3pqr

     putting values in above equation we get

      prq(3abc) - abc(3pqr) = 0